OpenStudy (anonymous):

9x - y = -2 Part A: Explain how you will show all of the solutions that satisfy this equation. (4 points) Part B: Determine three different solutions for this equation. (4 points) Part C: Write an equation that can be paired with the given equation in order to form a system of equations that is inconsistent. (2 points) can someone help me solve this please?

OpenStudy (aum):

9x - y = -2 solve for y. Add y: 9x = -2 + y add 2: 9x + 2 = y or y = 9x + 2 ---- (1) You can put any value of x in equation (1) and calculate y and those x and y values will satisfy the given equation 9x - y = -2. Since x can be any value and we can always calculate the corresponding y value from equation (1), there are infinite number of solutions to the equation 9x - y = -2.

OpenStudy (anonymous):

oh ok thanks

OpenStudy (aum):

For part B), pick any three values for x, put it in equation (1) and find the corresponding y. For example, if x = 2, then from equation (1): y = 9x + 2 = 9(2) + 2 = 18+2 = 20 Therefore, x = 2 and y = 20 is one solution. You can pick two more values for x and find the corresponding y values.

OpenStudy (anonymous):

ok I can do that thanks

OpenStudy (aum):

For part C) we need to come up with an equation that when paired with the given equation 9x - y = -2 becomes inconsistent. That can be easily done. For example, 9x - y = 5 is an equation that when paired up with 9x - y = -2 will be inconsistent because there is no x,y values that can both be 5 and -2 at the same time. Therefore, these two equations 9x - y = -2 and 9x - y = 5 will be inconsistent and there will be no solution to those two equations because no matter what value you give for x and y, the left hand side of both equations will be identical whereas the left hand side is -2 in the first equation and +5 in the second equation which cannot both be true.

OpenStudy (anonymous):

ok thank you

OpenStudy (aum):

yw. I noticed a couple of typos above which I will correct below: For part C) we need to come up with an equation that when paired with the given equation 9x - y = -2 becomes inconsistent. That can be easily done. For example, 9x - y = 5 is an equation that when paired up with 9x - y = -2 will be inconsistent because there is no x,y values that can simultaneously be 5 and -2 to satisfy both equations. Therefore, these two equations 9x - y = -2 and 9x - y = 5 will be inconsistent and there will be no solution to these pair of equations because no matter what value you give for x and y, the left hand side of both equations will be identical whereas the right hand side is -2 in the first equation and +5 in the second equation which cannot both be true.

OpenStudy (anonymous):

thanks